OpenStudy (anonymous):

A set of water waves travels at 10.0 m/s, and 5.0 waves pass you in 2.0 s. What is the wavelength of the waves?

OpenStudy (anonymous):

I know the answer is 4.0m i just need it explained to me how that answer came about

OpenStudy (anonymous):

v = a.f v = velocity a= the wavelength f = frequency aqual she spreads

OpenStudy (anonymous):

How would the equation look

OpenStudy (anonymous):

finding the frequency 5 --- 2 x --- 1 2x = 5 x = 5/2 x = 2.5 10 = a.f 10 = a.2,5 a = 10/2,5 a = 4

OpenStudy (anonymous):

ok thank you that helps